burning sodium? temperature increases by 2.03 °C. how much energy (in kJ/mol) is released by 3.03w 2.3 7,0125灯 5645 M.w 23 11. 12.0 g of water is at 57.0 ℃. If 4.00 g of gold at 18.2 °C is placed in the calorimeter, what is the final temperature of the water in the calorimeter? (The specific heat of gold is 0.128 J/g 12. A student wants to determine the identity of an unknown metal. The students heat the 25.0g chunk of metal to 122.8. °C, then drops it into 50.0g of water at 34.7 C. After waiting, the final temperature of the water is 38.5 C. (Use 4.184 J/g C as the specific heat of water.) From the following metals, what is the best guess for the identity of the unknown metal? m (snecific heat capacity of 0.921 J/g C)
Answer #1

Ans 11 :

The heat lost by water will be gained by gold

Heat = mass of substance x specific heat x change in temperature

Let the final temperature be T

specific heat of water = 4.186 J/goC

12.0 x 4.186 x (57.0 - T) = 4.00 x 0.128 x (T - 18.2)

2863.224 - 50.232 T = 0.512 T - 9.3184

2863.224 + 9.3184 = 0.512 T + 50.232 T

2872.5424 = 50.744 T

T = 56.6

so the final temperature of the water in calorimeter will be 56.6oC